Transverse waves travel through a string where the tension equals 14.58 N with a speed of 20.61 m/s. What tension would be required for a wave speed of 25.36 m/s?



Round your answer to 2 decimal places.


Do not forget the unit

1
Expert's answer
2022-05-11T13:52:53-0400
"v=\sqrt{F/\mu}"

"F_2=F_1\frac{v_1^2}{v_2^2}"

"F_2=14.58*\frac{25.36^2}{20.61^2}=22.1\:\rm N"